OpenStudy (anonymous):

Please help! Will give medal+fan! 1. What is the value of x? 4(6 + 7) = 4 • 6 + x • 7 2. Which expressions show how 5 • 37 can be rewritten using the distributive property? Choose all answers that are correct. A. 5 • 30 + 7 B. 5 • 30 + 5 • 7 C. 5 • 40 – 5 • 7 D. 5 • 40 – 5 • 3 3. Simplify: 5x3 – 2x2 + x + 4x3 A. 8x3 B. 9x6– 2x2+ x C. 9x3– 2x2+ x D. 5x3– 2x2+ x + 4x3 (the expression cannot be simplified) Thanks :D

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

@Improvement 8th grade

OpenStudy (anonymous):

this is for 5th graders

OpenStudy (anonymous):

5th grade students

OpenStudy (acxbox22):

@Improvement , stop bragging and if you know how to answer it, answer it

OpenStudy (mokeira):

\[4(6+7)=4.6+x.7\] This is the same as \[4(6+7)=(4\times6)+(x \times7)\] \[4(13)=24+7x\] \[52=24+7x\] \[52-24=7x\] \[28=7x\] \[4=x \]

OpenStudy (mokeira):

I dont know number 2

OpenStudy (mokeira):

\[5\times3-2\times2+x+4\times3\] \[15-4+x+12\] \[15+12-4+x\] \[23+x \] which is similar to choice C